OpenStudy (anonymous):

For the following problems show how the problem is set up, the work to solve the problem, along with the solution. Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart? I know the answer is 3 hours, but Idk how I would make a formula to find it and yeah

OpenStudy (anonymous):

@lalaly Could you help me with this quick question?

OpenStudy (anonymous):

I know the answer, just don't know how I would get a formula that would give me the answer.

OpenStudy (anonymous):

And that's a problem cause I need to show my work.

OpenStudy (lalaly):

\[Velocity = \frac{displacement}{time}\]\[s =Velocity \times time\] \[45=(10t+5t)\]\[45=15t\]\[t=3hrs\]

OpenStudy (anonymous):

Tyvm @lalaly

OpenStudy (lalaly):

your welcome:D

OpenStudy (anonymous):

Gonna write a testimonial right now for you :)

OpenStudy (lalaly):

haha thats Sweet:D Can't wait

OpenStudy (anonymous):

Can't write now though but later I will :) and how do I write one? NEver written one before..

OpenStudy (lalaly):

lol the thought of it is enough for me :D .. click on the name u will get something like this it says click here to write a tesetimonial

OpenStudy (anonymous):

K Thx :)

OpenStudy (lalaly):

:D

OpenStudy (anonymous):

Cute picture by the way lala